OpenStudy (anonymous):

Find the area of a square whose side is 4+3square root of 6ft

OpenStudy (angela210793):

is it \[\sqrt[3]{6}?\]

OpenStudy (anonymous):

\[(4+3\sqrt{6})^2\]

OpenStudy (angela210793):

or \[3\sqrt{6}?\]

OpenStudy (anonymous):

\[16+24\sqrt{6}+54\] \[70+24\sqrt{6}\]